OpenStudy (anonymous):

Can someone explain to me how to do this problem: Find an equation of the line that is tangent to the graph of f and parallel to the given line. f(x) = 2x^2 4x+y+3=0

OpenStudy (anonymous):

Lets work on this question together

OpenStudy (anonymous):

Now to find the tangent line to the graph of f we need to find the derivative of f. For this particular function you can the power rule. Can you find the derivative?

OpenStudy (anonymous):

\[f \prime (x) = 4x\]

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

now, since we are looking for tangent line that is parallel to the given line, we have to find the slope of that given line, can you find the slope of the given line?

OpenStudy (anonymous):

4x

OpenStudy (anonymous):

are you sure?? check again

OpenStudy (anonymous):

?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

Well, thats still not quite right: here it is, you have 4x+y+3=0, now lets solve for y to get this into y=mx+b form. Subtract 4x from both sides and you will get: y+3=-4x. Now subtract 3 from both sides and you will get: y=-4x-3. Thus our slope is -4.

OpenStudy (anonymous):

Now since the slope of this is -4, we can set it equal to f'(x) because parrallel lines mean that the slope is equal.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so, we have : 4x=-4

OpenStudy (anonymous):

can you solve for x?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

now, lets find the y coordinate, by plugging negative 1 back into the orginal equation: can you evluate f(-1): 2x^2

OpenStudy (anonymous):

-2x^2

OpenStudy (anonymous):

well, remeber f(-1) means that you replace all the x with -1, So we have: f(-1)=2(-1)^2. Can you evluate that?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

oh yeah, I forgot about that. In that case its 4.

OpenStudy (anonymous):

order of operations: exponents first, then multiplication

OpenStudy (anonymous):

Thus: 2*(-1)^2

OpenStudy (anonymous):

square -1 first, then multiply by 2

OpenStudy (anonymous):

2 then

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

now we have what we need, we have the y-coordinate: 2 x:coordinate: -1 and the slope: -4

OpenStudy (anonymous):

now with this information we can use the point slope form to get the equation of the tangent line to f, can you do that?

OpenStudy (anonymous):

\[2 = -4(1)+b\]\[b=6, y = -4x+6\]

OpenStudy (anonymous):

y-2=-4(x-(-1))

OpenStudy (anonymous):

y-2=-4(x+1)

OpenStudy (anonymous):

y-2=-4x-4

OpenStudy (anonymous):

oh, so I have to do it that way then. Okay...

OpenStudy (anonymous):

y=-4x-2

OpenStudy (anonymous):

and that is your final answer

OpenStudy (anonymous):

Yes, I'm trying to find the derivative of this function using the chain rule and the product rule. \[y = x \sqrt{1-x^{2}}\]

OpenStudy (anonymous):

okay then, so the product rule goes like F'*S+S'*F. F=first term S=second term

OpenStudy (anonymous):

so lets do this: what is the derivative of the first term x?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

right, now what is the derivative of the second term, you will have to use the chain rule here. Hint: write the second term as (1-x^2)^1/2

OpenStudy (anonymous):

\[1/2(1-x)^{-1/2}(-2x)\]

OpenStudy (anonymous):

it should be 1-x^2 inside the first parentheses right?

OpenStudy (anonymous):

should be: \[1/2(1-x^2)^-1/2*(-2x)\]

OpenStudy (anonymous):

4

OpenStudy (anonymous):

yes, sorry bout that

OpenStudy (anonymous):

okay, now lets finish it up by using the product rule

OpenStudy (anonymous):

we have : (1)(1-x^2)^1/2+ 1/2(1-x^2)^-1/2(-2x)(x)

OpenStudy (anonymous):

we can clean this up: (1-x^2)^1/2+(-2x^2)./2(1-x^2)^1/2

OpenStudy (anonymous):

now all we do is get lcd. Can you do that?

OpenStudy (anonymous):

Does that last part say 2(1-x^2)^1/2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So does that whole thing say: \[(1-x ^{2})^{1/2}+(-2x ^{2})/2(1-x ^{2})^{1/2}\] if so how did you get \[2(1-x ^{2})^{1/2}\]

OpenStudy (anonymous):

yes thats righ

OpenStudy (anonymous):

look at the post right underneath where i write"okay, now lets finish it up by using the product rule"

OpenStudy (anonymous):

basically, first i factores out the 1/2, then i just mulitplied (-2x)*(x). Next, i simply made (1-x^2)^-1/2 into (1)/(1-x^2)^1/2. Finally i multipled that by (-2x^2)/2

OpenStudy (anonymous):

(-2x^2)/2 comes from (1/2)*(-2x^2)

OpenStudy (anonymous):

okay I think I get it now.

OpenStudy (anonymous):

alright:)

OpenStudy (anonymous):

So -2x^2/2 is the answer?

OpenStudy (anonymous):

rmeber you need to get lcd to fully simplify

OpenStudy (anonymous):

so not the answer not -2x^2/2

OpenStudy (anonymous):

here is what we have: \[(1-x^2)^{1/2}-((2x^2)/(2(1-x^2)^{1/2})\]

OpenStudy (anonymous):

now you have to find the lcd and simplify

OpenStudy (anonymous):

so (1-x^2) is the lcd...

OpenStudy (anonymous):

2(1-x^2)^1/2 is the lcd

OpenStudy (anonymous):

Is this right so far: \[2(1-x ^{2})^{1/2}/2(1-x ^{2})^{1/2} \times (1-x ^{2})^{1/2}/2(1-x ^{2})^{1/2} - 2x ^{2}/2(1-x ^{2)^{1/2}}\] Everything on the left side cancels out except for the 2 So your left with 1/2. If you flip that and bring the 2 on top you have: \[2 \times (1) - 2x ^{2} /2(1-x ^{2})^{1/2}\]

OpenStudy (anonymous):

the 2;s in the numerator and denominator cancel out so you are left with on (1-2x^2)/(1-x^2)^1/2

OpenStudy (anonymous):

Then if I change that denominator to \[\sqrt{1-x ^{2}}\] my final answer, is then \[1-2x ^{2}/\sqrt{1-x ^{2}}\]

OpenStudy (anonymous):

that is correct